OpenStudy (solomonzelman):

My teacher did one example in class. I'll do a problem, and anyone is invited....If I'm wrong reply or message, but please don't do my prob. for me.

OpenStudy (solomonzelman):

I'll be posting part by part....

OpenStudy (solomonzelman):

I'll be posting part by part....

OpenStudy (solomonzelman):

The position of a particle is given by the equation \(\normalsize\color{blue}{ s=f(t)=t^3-6t^2+9t }\) where t is measured in seconds and s in meters. a) Find the velocity at time t. b) What is the velocity after 2 s? After 4 s? c) When is the particle at rest? d) When is the particle moving forward (that is in the positive direction) ? e) Draw a diagram to represent the motion of the particle. f) Find the total distance traveled by the particle during the 1st 5 seconds. g) Find the acceleration at time t and after 4 s.

OpenStudy (solomonzelman):

for part a, the velocity function, is the derivative of the position function. \(\normalsize\color{blue}{ v(t)=3t^2-12t+9 }\) ( v(t) is the new velocity function)

OpenStudy (solomonzelman):

for part b, the velocity after 2 s, means the instantaneous velocity, when t=2. So, \(\normalsize\color{blue}{ v(2)=3(2)^2-12(2)+9=12-24+9=-3 }\) (This is ds/dt at t=2.) SO after 2 s, it is \(\normalsize\color{blue}{ -3~m/s }\) The velocity after 4 s, is ds/dt at t=4, so \(\normalsize\color{blue}{ v(4)=3(4)^2-12(4)+9=48-48+9=9 }\) SO after 4 s, it is \(\normalsize\color{blue}{ 9~m/s }\).

OpenStudy (solomonzelman):

For part c, we know that the particle is at rest when t=0. so we set the velocity function equal to zero. \(\normalsize\color{blue}{ v(t)=3t^2-12t+9=0 }\) Solving for t, \(\normalsize\color{blue}{ 3t^2-12t+9=0 }\) \(\normalsize\color{blue}{ 3(t^2-4t+3)=0 }\) \(\normalsize\color{blue}{ 3(t-1)(t-3)=0 }\) So the particle is at rest after 3 s, and 1s.

OpenStudy (solomonzelman):

Part d. The particle is moving in a positive direction when \(\normalsize\color{blue}{ v(t)>0}\) , and that is: \(\normalsize\color{blue}{ 3(t-1)(t-3)>0 }\) . the inequality is true when both factors are positive or when both are negative, and thus \(\normalsize\color{blue}{ t<1 }\) to make both factors negative. and \(\normalsize\color{blue}{ t>3 }\) to make both factors positive.

OpenStudy (solomonzelman):

For part e. schematic sketch of the motion of the particle.  this is the motion back and forth along the s-axis. (s plays the role of the x)